# Dissipation-Driven Topological Phase Transition in Quantum Open Systems Independent of system Hamiltonian

### v1

@@ Abstract
We investigate dissipation-driven topological phase transitions in one-dimensional quantum open systems governed by the Lindblad equation with linear dissipation operators, which ensure the density matrix retains its Gaussian form throughout the dynamics. By employing the modular Hamiltonian framework, we rigorously demonstrate that the $\rm{Z}_{2}$ topological invariant characterizing steady states in one-dimensional class D systems is exclusively dependent on the dissipation operators, rather than the system Hamiltonian. Through a quench protocol where the system evolves from the steady state of one Lindbladian to another, we reveal that topological transitions can occur at analytically predictable critical times, even when the initial and final steady states share identical topological indices. These transitions are shown, both theoretically and numerically, to depend solely on dissipation parameters. Entanglement spectrum analysis demonstrates bulk-edge correspondence in non-equilibrium density matrices via coexisting single-particle gap closures (periodic boundaries) and topologically protected zero modes (open boundaries), directly underpinning the detection of dissipation-induced topology in quantum simulators.
@@ I Introduction
The investigation of topological insulators and superconductors in closed quantum systems constitutes a well-established research paradigm within condensed matter physics[1, 2, 3]. Here, the topological properties of a Hamiltonian are intrinsically encoded in its eigenstates, exemplified by bulk-boundary correspondence and quantized invariants like Chern numbers derived from Berry phases. However, the situation becomes fundamentally different in open quantum systems governed by Lindblad dynamics[4, 5]. Here, the system’s evolution is described by a Liouvillian superoperator rather than a Hamiltonian, and quantum states must be treated as mixed density matrices. Two distinct paradigms emerge for characterizing topology in such dissipative settings. One focuses on the intrinsic topology of mixed states through their density matrix structure[6, 7, 8, 9, 10, 11, 12, 13, 14], while the other explores the non-Hermitian topology[15, 16, 17, 18, 19, 20, 21, 22] encoded in the Liouvillian superoperators. Recent theoretical advances have established the modular Hamiltonian formalism ($\hat{\rho}=e^{-\hat{G}}$) as a rigorous framework for characterizing topology in quadratic Lindbladian systems[23, 24, 25, 26], where $\hat{\rho}$ remains Gaussian and $\hat{G}$ stays quadratic. The Hermitian nature of $\hat{G}$ allows direct application of topological classification schemes of closed systems, provided its symmetry class remains preserved during dissipative evolution. This preservation imposes strict requirements for both Lindbladian operators $\{\hat{L}_{\mu}\}$ and the system Hamiltonian. There exists symmetry classification correspondence between Gaussian states and quadratic Lindbladians [27]. While existing studies have established that topological transitions necessarily occur at finite critical times $t_{0}$ when initial and steady states occupy distinct topological sectors, it remains unknown whether multiple topological transitions can occur even when initial and final states are in different topological sectors, or how to determine their exact occurrence times in symmetry-constrained open systems.
@@
Here, we study one-dimensional Lindblad systems with Majorana fermions, where the Hamiltonian takes a quadratic form and dissipation operators are linear in Majorana operators. The modular Hamiltonian of such systems inherently satisfies particle-hole symmetry, naturally placing it within the class D framework without time-reversal symmetry[3]. Crucially, We find the steady state depends exclusively on the dissipation operators rather than the system Hamiltonian. To probe topological transitions, we employ a quench protocol: the system is initialized in a Gaussian state and subsequently evolved under a Lindbladian $\mathcal{L}$. We demonstrate that even when initial and final states share the same $Z_{2}$ topological indices, the dissipative dynamics can drive multiple topological transitions at precisely predictable time points. These critical times are fully determined by the dissipation operators, revealing a complete decoupling from Hamiltonian-driven dynamics observed in closed systems.
@@ II Model
where we have defined the Majorana operators as $\hat{w}_{m,1}=\frac{1}{\sqrt{2}}\left(\hat{c}_{m}^{\dagger}+\hat{c}_{m}\right)$ and $\hat{w}_{m,2}=i\frac{1}{\sqrt{2}}\left(\hat{c}_{m}-\hat{c}_{m}^{\dagger}\right)$, and commutation relation is written as $\{\hat{w}_{m,s},\hat{w}_{n,s^{\prime}}\}=\delta_{m,n}\delta_{s,s^{\prime}}$. The linear Lindblad operator $\hat{L}_{\mu}$ in Eq.(1) is superposition of both annihilation operators and creation operators, which takes the form
@@
By substituting Eq. (6) into Eq. (1), we obtain the time-evolution equation of the correlation matrix $\Delta$ as
@@
The matices $K$, $X$ and $Y$ are all transformed into the quasimomentum space as $K_{s,s^{\prime}}(k)=\sum_{m,n}K_{m,s,n,s^{\prime}}e^{ikm}e^{-ikn}$, $X_{s,s^{\prime}}(k)=\sum_{m,n}X_{m,s,n,s^{\prime}}e^{ikm}e^{-ikn}$, $Y_{s,s^{\prime}}(k)=\sum_{m,n}Y_{m,s,n,s^{\prime}}e^{ikm}e^{-ikn}$. The anti-symmetry of the matrices $K$, $Y$ and $\Delta$ and the symmetry of the matrix $X$ acts on quasimomentum space as $K(k)=-K^{T}(-k)$, $Y(k)=-Y^{T}(-k)$, $\Delta(k)=-\Delta^{T}(-k)$ and $X(k)=X^{T}(-k)$.
Since $\Delta$ is anti-symmetric, the modular Hamiltonian possesses an intrinsic particle-hole symmetry. In our context, we focus on D class which exhibits particle-hole symmetry but lacks chiral symmetry. The topological properties of one-dimensional class D systems can be characterized by a $\rm{Z}_{2}$ topological index $\nu$, which is calculated by Pfaffians of $\Delta$ as
@@
III Topological phase disgram of steady states
@@
Through vectorizing matrices $\Delta$ and $Y$ and expressing $H_{\text{eff}}$ in its eigenbasis $\{|\psi_{m}\rangle\}$ with eigenvalues , the steady-state correlation matrix of this equation transforms into
$\Delta_{s}(k)=4i\sum_{m,n=\pm}\frac{\langle\chi_{m}|Y(k)|\chi_{n}\rangle}{\lambda_{m}-\lambda_{n}^{*}}|\psi_{m}\rangle\langle\psi_{n}|,$ (15)
where $|\psi_{m}\rangle$ and $|\chi_{m}\rangle$ denote the right and left eigenvectors of $H_{\text{eff}}$ respectively and $\lambda_{m}$ ($m=\pm$) denotes the eigenvalues of $H_{\rm{eff}}$. Detailed deduction of Eq. (15) is shown in Appendix A. At high-symmetric points, we have $H_{\text{eff}}(k_{s})=2[K(k_{s})-iX(k_{s})]$, where $K(k_{s})$ is an antisymmetric Hermitian matrix and $X(k_{s})$ is a symmetric Hermitian matrix. Therefore, $H_{\text{eff}}(k_{s})$ could be written as
@@
${\rm{Pf}}[i\Delta_{s}(k_{s})]=\frac{1}{2}{\rm{Tr}}[\Delta(k_{s})\sigma_{y}]=\sum_{m,n}\frac{2iy_{k_{s}}f_{m,n}}{\lambda_{m}-\lambda_{n}^{*}},$ (19)
@@
$f_{m,n}=\langle\chi_{m}|\sigma_{y}|\chi_{n}\rangle\langle\psi_{n}|\sigma_{y}|\psi_{m}\rangle.$ (20)
Given Eq. (16), it follows directly that $H^{*}_{\rm{eff}}(k_{s})=-H_{\rm{eff}}(k_{s})$. Thus we find $f_{+,+}=f_{-,-}=1(0)$ and $f_{+,-}=f_{-,+}=0(1)$ for the PT-symmetry-preserved (broken) regime. We then substitute the respective overlap factors $f_{\pm,\pm}$ of both regimes into the Pfaffian’s expression Eq. (19). Since $\lambda_{+}(k_{s})-\lambda_{+(-)}^{*}(k_{s})=2ih_{0}$ holds for the PT-symmetry-preserved (broken) regime, this substitution process yields identical expressions for the Pfaffian of the correlation matrix in both regimes as
@@
$\hat{L}_{n}=u_{1}\hat{c}_{n}+u_{2}\hat{c}_{n+1}+v_{1}\hat{c}_{n}^{\dagger}+v_{2}\hat{c}_{n+1}^{\dagger},$ (27)
@@
Figure 1: Topological phase diagram of Lindblad steady state. Different colors denote different $(M_{0},M_{\pi})$. Red area represents $(M_{0},M_{\pi})=(-1,-1)$, blue area represents $(M_{0},M_{\pi})=(-1,1)$, green area represents $(M_{0},M_{\pi})=(1,-1)$ and white area represents $(M_{0},M_{\pi})=(1,1)$. Also, we have taken $v_{2}/v_{1}=-2$. Note that the system Hamiltonian parameters $J$, $\delta$ and $\mu$ has no influence on the topological properties of steady state.
@@ IV Topological phase transition in the dissipative dynamics
$\Delta(k_{s},t)-\Delta_{s,f}(k_{s})=\sum_{m,n}\exp[-i(\lambda_{m,f}-\lambda_{n,f}^{*})t]$
$\times\langle\chi_{m,f}|[\Delta_{s,i}(k_{s})-\Delta_{s,f}(k_{s})]|\chi_{n,f}\rangle|\psi_{m,f}\rangle\langle\psi_{n,f}|,$ (30)
where $\lambda_{m,f}$ is the eigenvalue of the effective non-Hermitian Hamiltonian in the post-quench Lindbladian $\mathcal{L}_{f}$, with $|\psi_{m,f}\rangle$ and $|\chi_{m,f}\rangle$ being its respective right and left eigenvectors. According to Eq. (11), the topological invariant of the evolving density state is directly calculated from the Pfaffians of correlation matrix
@@
${\rm Pf}[i\Delta(k_{s},t)]=$ ${\rm Pf}[i\Delta_{s,f}(k_{s})]+\frac{1}{2}\sum_{m,n}\exp[-i(\lambda_{m,f}-\lambda_{n,f}^{*})t]$
$\times{\rm\{{Pf}}[i\Delta_{s,i}(k_{s})]-{\rm{Pf}}[i\Delta_{s,f}(k_{s})]\}f_{m,n}.$ (32)
For the PT-symmetry-preserved regime, we only need to consider the $m=n$ case which yields $\lambda_{m}-\lambda_{n}^{*}=2ih_{0,f}$. And for the PT-symmetry-broken regime, we only consider the $m\neq n$ case also yields $\lambda_{m}-\lambda_{n}^{*}=2ih_{0,f}$. In both scenarios, Eq. (IV) is simplified to
@@
A dynamical topological phase transition takes place at points when topological index $\nu(t)=M_{0}M_{\pi}={\rm{Pf}}[i\Delta(0,t)]{\rm{Pf}}[i\Delta(\pi,t)]$ flips, which means either ${\rm{Pf}}[i\Delta(0,t)]$ or ${\rm{Pf}}[i\Delta(\pi,t)]$ crosses zero. The equation ${\rm{Pf}}[i\Delta(k_{s},t)]=0$ has at most one real solution, given by
@@
Figure 2: (a)(b)(c)Topological phase transition characterized by the modular Hamiltonian’s topological invariant $\nu(t)$ following a dissipative quench from Lindbladian steady state of $(M_{0},M_{\pi})=(1,-1)$ topological phase to those of $(M_{0},M_{\pi})=(1,1)$ phase. During the entire quench process, $v_{2}$ and $v_{1}$ remain constant, and the ratio $v_{2}/v_{1}$ is maintained at 2-2. Initial density matrix is the steady state of the Lindbladian with parameters $u_{1,i}/v_{1}=3$, $u_{2,i}/v_{1}=3$ and the post-quench Lindbladian parameters are taken as $u_{1,f}/v_{1}=2.5,u_{2,f}/v_{1}=-1$ for (a), $u_{1,f}/v_{1}=3.5,u_{2,f}/v_{1}=-1$ for (b), $u_{1,f}/v_{1}=4.5,u_{2,f}/v_{1}=-1$ for (c). (d)Topological phase transition time, prediected by $t_{p,0}$, varies with the post-quench parameter $u_{1,f}$.
Figure 3: (a)(b)(c)Topological phase transition characterized by the modular Hamiltonian’s topological invariant $\nu(t)$ following a dissipative quench from Lindbladian steady state of $(M_{0},M_{\pi})=(-1,-1)$ topological phase to those in $(M_{0},M_{\pi})=(1,1)$ phase. During the entire quench process, $v_{2}$ and $v_{1}$ remain constant, and the ratio $v_{2}/v_{1}$ is maintained at 2-2. Initial density matrix is the steady state of the Lindbladian with parameters $u_{1,i}/v_{1}=1$, $u_{2,i}/v_{1}=-1$ and the post-quench Lindbladian parameters are taken as $u_{1,f}/v_{1}=2.5,u_{2,f}/v_{1}=-1$ for (a), $u_{1,f}/v_{1}=3.5,u_{2,f}/v_{1}=-1$ for (b), $u_{1,f}/v_{1}=4.5,u_{2,f}/v_{1}=-1$ for (c). (d)Critical time of topological phase transition prediected by $t_{p,0}$ and $t_{p,\pi}$ vary with the post-quench parameter $u_{1,f}$ .

### v2

# Dissipation-driven topological phase transitions in open quantum systems independent of system Hamiltonian

@@ Abstract
We investigate dissipation-driven topological phase transitions in one-dimensional quantum open systems governed by the Lindblad equation with linear dissipation operators, which ensure the density matrix retains its Gaussian form throughout the dynamics. By employing the modular Hamiltonian framework, we rigorously demonstrate that the $\rm{Z}_{2}$ topological invariant characterizing steady states in one-dimensional class D systems is exclusively dependent on the dissipation operators, rather than the system Hamiltonian. Through a sudden quench protocol where the system evolves from the steady state of one Lindbladian to another, we reveal that topological transitions can occur at analytically predictable critical times, even when the initial and final steady states share identical topological indices. These transitions are shown, both analytically and numerically, to depend solely on dissipation parameters. Entanglement spectrum analysis demonstrates bulk-edge correspondence in non-equilibrium density matrices via coexisting single-particle gap closures (periodic boundaries) and topologically protected zero modes (open boundaries), directly underpinning the detection of dissipation-induced topology in quantum simulators.
@@ I Introduction
The investigation of topological insulators and superconductors in closed quantum systems constitutes a well-established research paradigm within condensed matter physics[19, 37, 13]. Here, the topological properties of a Hamiltonian are intrinsically encoded in its eigenstates, exemplified by bulk-boundary correspondence and quantized invariants like Chern numbers derived from Berry phases. However, the situation becomes fundamentally different in open quantum systems governed by Lindblad dynamics[28, 18]. Here, the system’s evolution is described by a Liouvillian superoperator rather than a Hamiltonian, and quantum states must be treated as mixed density matrices. Two distinct paradigms emerge for characterizing topology in such dissipative settings. One focuses on the intrinsic topology of mixed states through their density matrix structure[5, 15, 23, 41, 40, 12, 6, 43, 21], while the other explores the non-Hermitian topology[47, 46, 16, 24, 3, 27, 11, 34] encoded in the Liouvillian superoperators. In parallel, field-theoretic approaches have established a comprehensive classification and the bulk-boundary correspondence for Gaussian steady states across all symmetry classes and dimensions[22, 21], with extensions to interacting systems[20]. Recent theoretical advances have established the modular Hamiltonian formalism ($\hat{\rho}=e^{-\hat{G}}$) as a rigorous framework for characterizing topology in quadratic Lindbladian systems[2, 29, 42, 31], where $\hat{\rho}$ remains Gaussian and $\hat{G}$ stays quadratic. The Hermitian nature of $\hat{G}$ allows direct application of topological classification schemes of closed systems, provided its symmetry class remains preserved during dissipative evolution. This preservation imposes strict requirements for both Lindbladian operators $\{\hat{L}_{\mu}\}$ and the system Hamiltonian. There exists symmetry classification correspondence between Gaussian states and quadratic Lindbladians [30]. While existing studies have established that topological transitions necessarily occur at finite critical times $t_{0}$ when initial and steady states occupy distinct topological sectors, it remains unknown whether multiple topological transitions can occur even when initial and final states are in different topological sectors, or how to determine their exact occurrence times in symmetry-constrained open systems.
@@
Here, we study one-dimensional Lindblad systems with Majorana fermions, where the Hamiltonian takes a quadratic form and dissipation operators are linear in Majorana operators. The modular Hamiltonian of such systems inherently satisfies particle-hole symmetry, naturally placing it within the class D framework without time-reversal symmetry[13]. Crucially, we find the steady state depends exclusively on the dissipation operators rather than the system Hamiltonian.
To probe topological transitions, we employ a sudden quench protocol. The system is prepared in the steady state of an initial Lindbladian $\mathcal{L}_{i}$ and is then evolved under a constant post-quench Lindbladian $\mathcal{L}_{f}$. We demonstrate that even when initial and final states share the same $Z_{2}$ topological indices, the dissipative dynamics can drive multiple topological transitions at precisely predictable time points. These critical times are fully determined by the dissipation operators, revealing a complete decoupling from Hamiltonian-driven dynamics observed in closed systems.
@@ II Model
where $K$ is imaginary and antisymmetric matrix. We have defined the Majorana operators as $\hat{w}_{m,1}=\frac{1}{\sqrt{2}}\left(\hat{c}_{m}^{\dagger}+\hat{c}_{m}\right)$ and $\hat{w}_{m,2}=i\frac{1}{\sqrt{2}}\left(\hat{c}_{m}-\hat{c}_{m}^{\dagger}\right)$, and commutation relation is written as $\{\hat{w}_{m,s},\hat{w}_{n,s^{\prime}}\}=\delta_{m,n}\delta_{s,s^{\prime}}$. The linear Lindblad operator $\hat{L}_{\mu}$ in Eq.(1) is superposition of both annihilation operators and creation operators, which takes the form
@@
By substituting Eq. (6) into Eq. (1), we obtain the time-evolution equation of the correlation matrix $\Delta$ [48, 36, 7]as
@@
The matrices $K$, $X$ and $Y$ are all transformed into the quasimomentum space as $K_{s,s^{\prime}}(k)=\sum_{m,n}K_{m,s,n,s^{\prime}}e^{ikm}e^{-ikn}$, $X_{s,s^{\prime}}(k)=\sum_{m,n}X_{m,s,n,s^{\prime}}e^{ikm}e^{-ikn}$, $Y_{s,s^{\prime}}(k)=\sum_{m,n}Y_{m,s,n,s^{\prime}}e^{ikm}e^{-ikn}$. The anti-symmetry of the matrices $K$, $Y$ and $\Delta$ and the symmetry of the matrix $X$ acts on quasimomentum space as $K(k)=-K^{T}(-k)$, $Y(k)=-Y^{T}(-k)$, $\Delta(k)=-\Delta^{T}(-k)$ and $X(k)=X^{T}(-k)$.
Since $\Delta$ is anti-symmetric, the modular Hamiltonian possesses an intrinsic particle-hole symmetry. In our context, we focus on D class which exhibits particle-hole symmetry but lacks chiral symmetry. The topological properties of one-dimensional class D systems can be characterized by a $\rm{Z}_{2}$ topological index $\nu$, which is calculated by Pfaffians of $\Delta$ [25] as
@@
III Topological phase diagram of steady states
@@
Through vectorizing matrices $\Delta$ and $Y$ and expressing $H_{\text{eff}}$ in its eigenbasis $\{|\psi_{\alpha}\rangle\}$ with eigenvalues , the steady-state correlation matrix of this equation transforms into
$\Delta_{s}(k)=4i\sum_{\alpha,\beta=\pm}\frac{\langle\chi_{\alpha}|Y(k)|\chi_{\beta}\rangle}{\lambda_{\alpha}-\lambda_{\beta}^{*}}|\psi_{\alpha}\rangle\langle\psi_{\beta}|,$ (15)
where $|\psi_{\alpha}\rangle$ and $|\chi_{\alpha}\rangle$ denote the right and left eigenvectors of $H_{\text{eff}}$ respectively and $\lambda_{\alpha}$ ($\alpha=\pm$) denotes the eigenvalues of $H_{\rm{eff}}$. Detailed deduction of Eq. (15) is shown in Appendix A. Note that the derivation of Eq.(15) requires the prerequisite that $\lambda_{\alpha}-\lambda_{\beta}^{*}\neq 0$ for all eigenvalues, a condition which is equivalent to a non-vanishing dissipative gap[36]. This prerequisite is assumed to hold throughout our work.
At high-symmetric points, we have $H_{\text{eff}}(k_{s})=2[K(k_{s})-iX(k_{s})]$, where $K(k_{s})$ is an antisymmetric Hermitian matrix and $X(k_{s})$ is a symmetric Hermitian matrix. Therefore, $H_{\text{eff}}(k_{s})$ could be written as
@@
${\rm{Pf}}[i\Delta_{s}(k_{s})]=\frac{1}{2}{\rm{Tr}}[\Delta(k_{s})\sigma_{y}]=\sum_{\alpha,\beta}\frac{2iy_{k_{s}}f_{\alpha,\beta}}{\lambda_{\alpha}-\lambda_{\beta}^{*}},$ (19)
@@
$f_{\alpha,\beta}=\langle\chi_{\alpha}|\sigma_{y}|\chi_{\beta}\rangle\langle\psi_{\beta}|\sigma_{y}|\psi_{\alpha}\rangle.$ (20)
Given Eq. (16), it follows directly that $H^{*}_{\rm{eff}}(k_{s})=-H_{\rm{eff}}(k_{s})$. Thus we find $f_{+,+}=f_{-,-}=1(0)$ and $f_{+,-}=f_{-,+}=0(1)$ for the PT-symmetry-preserved (broken) regime. We then substitute the respective overlap factors $f_{\pm,\pm}$ of both regimes into the Pfaffian’s expression Eq. (19). Since $\lambda_{\alpha}(k_{s})-\lambda_{\beta}^{*}(k_{s})=2ih_{0}$ when $\alpha=\beta$ ($\alpha\neq\beta$) holds for the PT-symmetry-preserved (broken) regime this substitution process yields identical expressions for the Pfaffian of the correlation matrix in both regimes as
@@
$\hat{L}_{\mu}=\sum_{n}\sqrt{\gamma}[(u_{1}\delta_{\mu,n}+u_{2}\delta_{\mu+1,n})\hat{c}_{n}+(v_{1}\delta_{\mu,n}+v_{2}\delta_{\mu+1,n})\hat{c}_{n}^{\dagger}],$ (27)
@@
Figure 1: Topological phase diagram of Lindblad steady state. Different colors denote different $(M_{0},M_{\pi})$. Red area represents $(M_{0},M_{\pi})=(-1,-1)$, blue area represents $(M_{0},M_{\pi})=(-1,1)$, green area represents $(M_{0},M_{\pi})=(1,-1)$ and white area represents $(M_{0},M_{\pi})=(1,1)$. Also, we have taken $v_{2}/v_{1}=-2$, and the dimensionless parameter $\gamma=0.1$. Note that the system Hamiltonian parameters have no influence on the topological properties of the steady state, although the validity of this phase diagram requires that the Hamiltonian is chosen to ensure a non-vanishing dissipative gap.
@@ IV Topological phase transition in the dissipative dynamics
$\Delta(k_{s},t)-\Delta_{s,f}(k_{s})=\sum_{\alpha,\beta}\exp[-i(\lambda_{\alpha,f}-\lambda_{\beta,f}^{*})t]$
$\times\langle\chi_{\alpha,f}|[\Delta_{s,i}(k_{s})-\Delta_{s,f}(k_{s})]|\chi_{\beta,f}\rangle|\psi_{\alpha,f}\rangle\langle\psi_{\beta,f}|,$ (30)
where $\lambda_{\alpha,f}$ is the eigenvalue of the effective non-Hermitian Hamiltonian in the post-quench Lindbladian $\mathcal{L}_{f}$, with $|\psi_{\alpha,f}\rangle$ and $|\chi_{\alpha,f}\rangle$ being its respective right and left eigenvectors. According to Eq. (11), the topological invariant of the evolving density state is directly calculated from the Pfaffians of correlation matrix
@@
${\rm Pf}[i\Delta(k_{s},t)]=$ ${\rm Pf}[i\Delta_{s,f}(k_{s})]+\frac{1}{2}\sum_{\alpha,\beta}\exp[-i(\lambda_{\alpha,f}-\lambda_{\beta,f}^{*})t]$
$\times{\rm\{{Pf}}[i\Delta_{s,i}(k_{s})]-{\rm{Pf}}[i\Delta_{s,f}(k_{s})]\}f_{\alpha,\beta}.$ (32)
For the PT-symmetry-preserved regime, we only need to consider the $\alpha=\beta$ case which yields $\lambda_{\alpha}-\lambda_{\beta}^{*}=2ih_{0,f}$. And for the PT-symmetry-broken regime, we only consider the $\alpha\neq\beta$ case also yields $\lambda_{\alpha}-\lambda_{\beta}^{*}=2ih_{0,f}$. In both scenarios, Eq. (IV) is simplified to
@@
A dynamical topological phase transition takes place at points when topological index $\nu(t)=M_{0}M_{\pi}={\rm{sgn}}\{{{\rm{Pf}}[i\Delta(0,t)]{\rm{Pf}}[i\Delta(\pi,t)]}\}$ flips, which means either ${\rm{Pf}}[i\Delta(0,t)]$ or ${\rm{Pf}}[i\Delta(\pi,t)]$ crosses zero. The equation ${\rm{Pf}}[i\Delta(k_{s},t)]=0$ has at most one real solution, given by
@@
Figure 2: (a)(b)(c)Topological phase transition characterized by the modular Hamiltonian’s topological invariant $\nu(t)$ following a dissipative quench from Lindbladian steady state of $(M_{0},M_{\pi})=(1,-1)$ topological phase to those of $(M_{0},M_{\pi})=(1,1)$ phase. During the entire quench process, $v_{2}$ and $v_{1}$ remain constant, the ratio $v_{2}/v_{1}$ is maintained at 2-2, and the dimensionless parameter $\gamma$ takes the value $0.1$. Initial density matrix is the steady state of the Lindbladian with parameters $u_{1,i}/v_{1}=3$, $u_{2,i}/v_{1}=3$ and the post-quench Lindbladian parameters are taken as $u_{1,f}/v_{1}=2.5,u_{2,f}/v_{1}=-1$ for (a), $u_{1,f}/v_{1}=3.5,u_{2,f}/v_{1}=-1$ for (b), $u_{1,f}/v_{1}=4.5,u_{2,f}/v_{1}=-1$ for (c). (d)Topological phase transition time, prediected by $t_{p,0}$, varies with the post-quench parameter $u_{1,f}$.
Figure 3: (a)(b)(c)Topological phase transition characterized by the modular Hamiltonian’s topological invariant $\nu(t)$ following a dissipative quench from Lindbladian steady state of $(M_{0},M_{\pi})=(-1,-1)$ topological phase to those in $(M_{0},M_{\pi})=(1,1)$ phase. During the entire quench process, the ratio $v_{2}/v_{1}$ is maintained at 2-2, and the dimensionless parameter $\gamma$ takes the value $0.1$. Initial density matrix is the steady state of the Lindbladian with parameters $u_{1,i}/v_{1}=1$, $u_{2,i}/v_{1}=-1$ and the post-quench Lindbladian parameters are taken as $u_{1,f}/v_{1}=2.5,u_{2,f}/v_{1}=-1$ for (a), $u_{1,f}/v_{1}=3.5,u_{2,f}/v_{1}=-1$ for (b), $u_{1,f}/v_{1}=4.5,u_{2,f}/v_{1}=-1$ for (c). (d)Critical time of topological phase transition prediected by $t_{p,0}$ and $t_{p,\pi}$ vary with the post-quench parameter $u_{1,f}$ .